OpenStudy (anonymous):

Integral ((x)/(x^(2)+1)

OpenStudy (zepp):

Use u-substitution to get the answer.

OpenStudy (zepp):

\[\large \int \frac{x}{x^2+1}dx\]Let u = x^2+1, then du=\(2xdx\); \(\frac{1}{2}du=xdx\)\[\large \frac{1}{2}\int\frac{1}{u}du\]

OpenStudy (anonymous):

The answer i got is (1/2)ln(x^(2)+1) + C but for some reason its not right.

OpenStudy (zepp):

That looks correct for me.\[\large =\frac{1}{2}ln|x^2+1|+C\] I used absolute value since that's how I learned that.

OpenStudy (anonymous):

alright thank you

OpenStudy (zepp):

np.